Actions
20160503 » History » Revision 16
« Previous |
Revision 16/17
(diff)
| Next »
Hodgson, Paul, 04 May 2016 10:05
Ops/Comm Scheduling meeting¶
Location : CR11 in R3 @ 15:30 pm on Tuesday 3rd
Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402
Agenda¶
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands
- Move the tracker CAENbus module higher up the chain (Craig)
- Get an estimate for how long it takes to wire in the chiller in (Andy)
- DC report
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6185/2016-05-03-MICE-Hall-Plan.xlsx
- QP/QD work : http://micewww.pp.rl.ac.uk/attachments/6181/MICE_Electrical_Work_List_25APR16.docx
- Water system work plan
- Bellows installation and FC cooldown
- Tracker hall probe installation
- AOB
- Prescaled trigger discussion (Online group). If you are interested, please indicate your availability in the doodle:
http://doodle.com/poll/wnkk4aevq938keb8
- Prescaled trigger discussion (Online group). If you are interested, please indicate your availability in the doodle:
- DONM : 10/3/2016
Minutes¶
- Present : Yordan, Paul H, Paul S, Mike C, Vicky B, Josef F, Colin W, Andy N, Craig M, Steve B, Pierrick H, Ken L
- Actions from last meeting
- Generate appropriate table to store absorber pressure and material information (Janusz) : (almost done)
- DS readback current requires calibration (DL) : stands until close to powerup
- Issues arising from last data taking run need to be identified and actioned (Boyd) : Stands
- Get an estimate for how long it takes to wire in the chiller in (Andy) : Done - about 3-4 days.
- DC report
- Focus coil won't go back in since the bellows have been switched around and now the screw hole tolerances no longer match.
- When John gets back from holiday the FC will be pulled out and the bellows swapped
- Daresbury work went well. An ongoing work plan is shown below
- US team came out with a program of work, some of which can be done but because the racks are not fully commissioned the
full program of work can't be completed. - This week HPC have started to work on the compressor for the DS. They are concerned about the temperature of the air end - it's
too hot to touch. They plan to replace this in August. Colin:Can we just buy a new one rather than replace in August? Mike: What we
get is effectively new but we should put an order in. Colin:we should do a comprehensive job on fixing the compressor so need to
understand this issue.- Action Item: Vicky will investigate the state of the refurbished air end to see what we get for our money.
- Daresbury will allow the US to run into a short this week.
- Alan B : the 20 psi main relief valve for SSU needs replacing. Alan has a replacement but it needs installation. This can happen
when John is back next week.
- Work plan for the next week : http://micewww.pp.rl.ac.uk/attachments/6180/2016-05-02-Workplan.pdf
- Henry: Water plan goes into period when the DSA is closed. If we turn the water back on we may be in a situation where the
water pipes to the conventional magnets could blow. If the DSA is closed we can't get to it. Colin: Good point. Henry: We also
need the full ISIS running schedule spreadsheet. Colin:I've got something - I'll send it to you. - Action Item: Colin will put the ISIS dates for run-ups and machine work into his schedules to better reflect available access time.
- Henry: Water plan goes into period when the DSA is closed. If we turn the water back on we may be in a situation where the
- QP/QD work
- Discussed during the DC report
- Water system work plan : Colin:There is no day-by-day plan, but the plan for what is being done is understood. The plan is not going
to (ha) plan due to delays in getting an external contractor.- Action Item: We will make some form of water system schedule available (Colin)
- Bellows installation and FC cooldown : Bellows discussed during DC report. We plan to pump the FC down ASAP and then, if happy,
start the cooldown. - Tracker hall probe installation: Will start tomorrow. Craig and Melissa have a plan which will get the losses down to an acceptable level. This will only take a day or so.
- AOB
- Prescaled trigger discussion: Yordan is organising a phone meeting to develop the prescaled trigger. For those interested there
is a Doodle poll for this meeting. - Henry: Still looking for people to train as BLOCs. BLOC training will happen in June. An invitation will be given at the VC.
- Henry: Water detection in the trench? Pierrick: we have water detection in the trench but it's in the gutter and so shows water
detection all the time. We need to come up with a proper plan.- Action Item: Henry, Craig, Paul and Pierrick to look into optimising this for once and for all.
- Henry: Can Paul lower the target frame and check that the laser light levels haven't changed. Paul: Yes.
- Prescaled trigger discussion: Yordan is organising a phone meeting to develop the prescaled trigger. For those interested there
Updated by Hodgson, Paul over 7 years ago · 16 revisions